OpenStudy (anonymous):

Which expression results from applying the Identity Property of Multiplication to 5u+5v+8v +v +10 to get an equivalent expression?

OpenStudy (anonymous):

what a weird question.

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

80 approaches. What do do when everybody is a superguru?

OpenStudy (anonymous):

it is clear that \[5u+5v+8v +v +10=5u+14v+10\] but i would say that it is a result of the distributive law, not of the identity property of multiplication

OpenStudy (anonymous):

Make an algebraist of u yet...

OpenStudy (anonymous):

i was going to retire at 1000 medals, but amistre's example kept me going.

OpenStudy (anonymous):

i long for the days when i was a humble master

OpenStudy (anonymous):

could you explain what here has to do with the identity property of multiplication? surely not writing \[v=1\times v\] and then adding?

OpenStudy (anonymous):

Ok i need The correct Answer iknow how to do this i just need the answer please!!!

OpenStudy (anonymous):

Multiplicative Property.

OpenStudy (anonymous):

it says "which expression"

OpenStudy (anonymous):

five u plus five v plus nine v plus ten five u plus v times the quantity five plus eight end quantity plus v plus ten five times the quantity u plus v end quantity plus eight v plus v plus ten five u plus thirteen v plus v plus ten These are the choices

OpenStudy (anonymous):

not "what property"

OpenStudy (anonymous):

Yours is the multiplicative property.

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

ok Thanks

OpenStudy (anonymous):

what a dumb quesion!

OpenStudy (anonymous):

:-)

OpenStudy (anonymous):

Collecting like terms is indeed based on the distributive property of the field. 2x + 2x = (2 + 2)x = 4x

OpenStudy (anonymous):

Don't u start:-)

OpenStudy (anonymous):

right. distributive law. but look at the question!

OpenStudy (anonymous):

I dont really get the question either sorry.

OpenStudy (anonymous):

must be some education major wants you to write \[5u+5v+8v+1\times v+10\] and then say wow look, \[8+1=9\] that is the only thing i can figure out

OpenStudy (anonymous):

maybe

OpenStudy (anonymous):

satellite73 Can i have your acount

OpenStudy (anonymous):

of course i could be wrong, but if you read the question i cannot figure out a different answer

OpenStudy (anonymous):

i have no account. i am a no account.

OpenStudy (anonymous):

What

OpenStudy (anonymous):

The identity property of multiplication, also called the multiplication property of one says that a number does not change when that number is multiplied by 1.

OpenStudy (anonymous):

Drum roll....

OpenStudy (anonymous):

There are additive and multiplicative identities. Additive identity is 0 x + 0 = x Multiplicative identity is 1 x1 = x

OpenStudy (anonymous):

What would we do without 0 and 1? There would be nowhere for all the analysts to go hang out.